OpenStudy (winsteria):

WILL MEDAL AND FAN What is the difference? 1/6 - 5/8

OpenStudy (chosenmatt):

-11/24

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }-\frac{ 5 }{ 8 }\] You need the denominators to match in order to subtract fractions so multiply the left side by 8 and the right side by 6. \[(\frac{ 1 }{ 6 }) *8 - (\frac{ 5 }{ 8 })*6\]\[\frac{ 8 }{ 48 }-\frac{ 30 }{ 48 }\] Only subtract the numerators, keep the denominators the same. \[-\frac{ 22 }{ 48 }\] Now reduce.\[-\frac{ 11 }{ 24 }\] so your answer is \[-\frac{ 11 }{ 24 }\]

OpenStudy (winsteria):

Thank you

OpenStudy (chosenmatt):

no problem